OpenStudy (anonymous):

3. There are 20 books on the summer reading list. In how many different ways can you choose one per week for six weeks? (1 point)

OpenStudy (anonymous):

Help!!

OpenStudy (anonymous):

20 choices for first book 19 choices for second 18 choices for third 17 choices for fourth 16 choices for fifth 15 choice for sixth counting principle says so multiply all these together

OpenStudy (anonymous):

your answer is \[20\times19\times 18\times 17\times 16\times 15\]

OpenStudy (anonymous):

@satellite73 I got D, 27,907,200...